A. A sample of oxygen gas at a pressure of 0.986 atm and a temperature of 23.3 °C, occupies a volume of 563 mL. If the gas is compressed at constant temperature until it’s pressure is 1.47 atm, the volume of the gas sample will be ______ mL.


1
Expert's answer
2022-03-30T02:09:01-0400

PV=nRT

nR=PV/T=0.986x563/T=555.1/T

1.47V=555.1T/T =555.1

V=555.1/1.47=377.6